OpenStudy (anonymous):

y varies inversely as three times x. When x = 6, y = 4. Find x when y = 8.

OpenStudy (anonymous):

I really dont understand what it means when Y varies inversely as three times X

OpenStudy (anonymous):

pretty sure it means \[y=\frac{k}{3x}\]

OpenStudy (anonymous):

Y=k/3x

OpenStudy (anonymous):

that is what it is.

OpenStudy (anonymous):

yes but I dont know how to solve it

OpenStudy (anonymous):

to solve for \(k\) put \[4=\frac{k}{3\times 6}\]

OpenStudy (anonymous):

that means \[4=\frac{k}{18}\iff k=18\times 4=72\]

OpenStudy (anonymous):

what is X when Y = 8

OpenStudy (anonymous):

since \[y=\frac{72}{3x}=\frac{24}{x}\] your job is to solve \[8=\frac{24}{x}\] for \(x\)

OpenStudy (anonymous):

you know how to do that?

OpenStudy (anonymous):

X=3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank You so much!

OpenStudy (anonymous):

yw